OpenStudy (anonymous):

write each number in scientific notation 4000

OpenStudy (unklerhaukus):

a number is in scientific notation when it is expressed as a digit (1-9) followed by a decimal, timesed by some power of ten like 2 = 2.0 x 10^0 356 = 3.56 x 10^2 0.0041 = 4.1 x 10^-3

OpenStudy (unklerhaukus):

7 000 000 = 7 x 10^6 0.00009981 = 9.981 x 10^-5

OpenStudy (unklerhaukus):

when the decimal point has moved to the left, the exponent of ten will be positive (and equal to the number of places it has moved) the the decimal point has moved to the right, the exponent of ten will be negative (and equal to the number of places the it has moved )

OpenStudy (unklerhaukus):

4 000. = 4.000 x 10 ^ ? how many places has the decimal point moved? will it be positive or negative?

OpenStudy (anonymous):

3

OpenStudy (unklerhaukus):

yes! , do you have more like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

6550000000

OpenStudy (unklerhaukus):

so whats the first step ?

OpenStudy (anonymous):

is it 655.0000000*10^7

OpenStudy (unklerhaukus):

that number is equal , so your on the right track, but to be scientific notation you need to only have one digit to the left of the decimal place

OpenStudy (anonymous):

65.50000000*10^7

OpenStudy (unklerhaukus):

that is still to many digits, on the left of the decimal place, also for that number to be equal you would have to add one to the index of the ten

OpenStudy (anonymous):

i give up

OpenStudy (unklerhaukus):

65.50000000*10^8 would be equal but you only want one digit to the left of the decimal place 6.55 * 10 ^ ?

OpenStudy (unklerhaukus):

(the zeros at on the right are insignificant so you dont have to write them)

OpenStudy (anonymous):

so your saying it will be 6.55*10^7

OpenStudy (unklerhaukus):

\[\large6\overbrace{~55\overbrace{0~000~000}^{7}}^?=655 \times 10^7=6.55\times10^?\]

OpenStudy (anonymous):

10

OpenStudy (unklerhaukus):

not quite try again \[\large6\overbrace{~55\overbrace{0~000~000.}^{7}}^?\\\large\qquad=655. \times 10^7\\\large\qquad=6.55\times10^?\]

OpenStudy (anonymous):

9

OpenStudy (unklerhaukus):

right

OpenStudy (anonymous):

my next one is 3940000

OpenStudy (unklerhaukus):

and what do think it is in scientific notation ? how many place does the point move

OpenStudy (anonymous):

i thing this one is 394*10^6

OpenStudy (unklerhaukus):

where'd the decimal point go?

OpenStudy (anonymous):

3.94

OpenStudy (unklerhaukus):

right, but now where'd the 10's go?

OpenStudy (unklerhaukus):

your almost there, just put it together

OpenStudy (anonymous):

3.94*10^6

OpenStudy (unklerhaukus):

YESS!!!

OpenStudy (anonymous):

next is 1220000

OpenStudy (anonymous):

12.2*10^5

OpenStudy (unklerhaukus):

well 1 220 000 is equal to 12.2*10^5, but remember you want only one digit to the left for scientific notation

OpenStudy (anonymous):

1.22*10^6

OpenStudy (unklerhaukus):

√ Right !

OpenStudy (anonymous):

next is 4.3

OpenStudy (unklerhaukus):

ok well you will need to have a x10^? on the end but what number should ? be has the decimal place moved>?

OpenStudy (anonymous):

ithnk it is 4.3*10^0

OpenStudy (unklerhaukus):

√√√ correct

OpenStudy (anonymous):

next is 11600

OpenStudy (anonymous):

i think it is 1.16*10^4

OpenStudy (unklerhaukus):

good work,

OpenStudy (anonymous):

next is 0.00000021

OpenStudy (unklerhaukus):

ok so this time we have to move the decimal place to the right, the power of 10 will be a negative number

OpenStudy (anonymous):

is it 2.1*10^-8

OpenStudy (unklerhaukus):

very close, count those places again i find it helps to rewrite the number with space 0.000 000 21

OpenStudy (anonymous):

-7

OpenStudy (unklerhaukus):

thats better!

OpenStudy (anonymous):

next one is 7150000

OpenStudy (anonymous):

is it 7.15*10^6

OpenStudy (unklerhaukus):

that is right

OpenStudy (anonymous):

now this is diffrent it says fix each number so that it is in scientific notation

OpenStudy (unklerhaukus):

ok

OpenStudy (anonymous):

first problem is 0.56*10^6

OpenStudy (unklerhaukus):

ah so its almost there, you just have to move the decimal place one step, is it positive or negative?

OpenStudy (anonymous):

so it will be 5.6*10^6

OpenStudy (unklerhaukus):

but you have to change the index of 10

OpenStudy (anonymous):

^7

OpenStudy (unklerhaukus):

the decimal place moved the right (not the left )

OpenStudy (anonymous):

.56

OpenStudy (unklerhaukus):

0.56*10^6 = 5.6*10^(6-1)

OpenStudy (anonymous):

next one is 320*10^3

OpenStudy (unklerhaukus):

this one the decimal place moves to the left, so ^(3+?)

OpenStudy (anonymous):

3

OpenStudy (unklerhaukus):

nope